Michelle has a party and invites 3 friends., each friend invites 3 other friends . each of these friends invites 3 more friends . Finally, each of these friends invites 3 more friends. each guest at the party brings 3 gifts with 3 bows on each gift. how many bows does michelle get at the party ?
Answer by god2012(113)

Solution :
Total number of guest at the party = 3*3*3*3 = 81
Total number of gifts = 81*3 = 243
Total number of bowls Michell gets = 243*3 = 729